CoolRekterizer:

3√8w^2 ... Solve this.

darkknight:

What are we trying to do

CoolRekterizer:

try to solve it

darkknight:

?

darkknight:

simplify?

CoolRekterizer:

\[\sqrt[3]{8w ^{2}}\]

CoolRekterizer:

yes

darkknight:

oh ok, so what you do is that can take the W^2 out of the square root. do you know what i mean?

darkknight:

like the w^2 when you square root is w

TheRobin:

2\[\sqrt[2]{8w ^{2}}\]

TheRobin:

sorry, the I meant to put a 3 with the radical

CoolRekterizer:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @TheRobin 2\[\sqrt[2]{8w ^{2}}\] \(\color{#0cbb34}{\text{End of Quote}}\) Its not sqaure root, its cubed

darkknight:

and simplifying the sqrt 8 u can rewrite as 4 times 2 under the square root then take the square root of 4 and get 2 times sqrt of 2

TheRobin:

Ik I meant to put three

CoolRekterizer:

i think i got it

CoolRekterizer:

\[2w ^{\frac{ 2 }{ 3 }}\]

darkknight:

no

CoolRekterizer:

i searched in mathway, its correct

TheRobin:

I also used mathway and I got different answer

darkknight:

so you have 3 times sqrt of 8w^2 right then you can take the w out like how I told u and have 3w times sqrt of 8. You can simplify the 8 into 4 times 2 and then take the 4 out getting 6w times sqrt 2

CoolRekterizer:

Use properties of exponents

justus:

your answer is right

CoolRekterizer:

my answer that i got was \[2w ^{\frac{ 2 }{ 3 }}\]

TheRobin:

whose?

CoolRekterizer:

ill even snip it

darkknight:

what math is this, algebra 1?

CoolRekterizer:

well for my brother

CoolRekterizer:

yea, he needs help

darkknight:

this is your brother's problem? can't u help him then?

CoolRekterizer:

im too lazy, and i practically dont understand

darkknight:

hold up, do you understand what I did for that problem?

CoolRekterizer:

1 attachment
darkknight:

all of them?

TheRobin:

Cool, that is the Rational Exponents, not simplified form, which one does it need?

darkknight:

already did 1 2: distribute the exponent that is to the power of 3 over all the individual terms, like the 8 is going to be 2, and the w^2 is going to be w^2/3 so the new eqn becomes 2(w^(2/3))

darkknight:

3: simplify the 8 like in the step above well it leads to the same answer as 2 xD

TheRobin:

1 attachment
darkknight:

4) is also 3 but the 8 is rewritten so it is also the same

darkknight:

oh wait a minute, these are all steps to the same problem

darkknight:

When you realize...

darkknight:

OMG, why??? you said it was 3 times 8w^2 first not cube root

darkknight:

Where do you see that?